Ask
your own question, for FREE!
Mathematics
7 Online
A coin is tossed 100 times resulting in 46 heads and 54 tails. The same coin is tossed 1000 times resulting in 495 heads and 505 tails. What is the theoretical probability of getting heads with this coin? A. 49.2% B. 49.5% C. 50% D. 46%
Still Need Help?
Join the QuestionCove community and study together with friends!
Out of 1100 tosses heads came up 45 + 495 = 541 times and tails came up 54 + 505 = 559 times. Therefore the experimental probability of getting heads is 541/1100. However if the coin is fair, the theoretical probability of getting heads is 1/2 = 0.5.
thank you
You're welcome :)
@dbrown13 Do you understand which answer option to choose now?
yes
Still Need Help?
Join the QuestionCove community and study together with friends!
Which one did you choose?
C.
Your choice is correct :)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
20 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal